# OmniLens: Towards Universal Lens Aberration Correction via LensLib-to-Specific Domain Adaptation

### v2

@@ 1 Introduction
Figure 1: OmniLens is a flexible solution to universal CAC. In real-world cases of unknown lens descriptions for the applied low-end lenses, the pre-trained model with manually collected LensLib [31] delivers poor performance and can hardly be further fine-tuned. In contrast, the automatically generated LensLib in OmniLens develops a more robust model, where the LensLib-to-specific domain adaptation can further enhance the results.
@@
Computational Aberration Correction (CAC), equipped with a post-image-restoration method to deal with the degradation (i.e., the optical degradation [5]) induced by residual optical aberrations of the target optical lens, is a fundamental and long-standing task [42, 27] in computational imaging. This technology is particularly highly demanded in mobile and wearable vision terminals for lightweight and high-quality photography [48, 62], where low-end lenses [18, 13] with simple structure and severe optical degradation are applied, as shown in Fig. 1 (a). Recent advances in CAC have centered around Deep Learning (DL) based methods [36, 4, 7] equipped with a powerful image restoration network [34, 3, 59] being trained on the corresponding data pairs under lens-specific optical degradation [4, 5]. However, these methods have invariably been tailored to a specific lens, revealing limited generalization ability to other different lens designs. In this way, the complex and time-consuming pipeline of lens description calibration, imaging simulation, data preparation, and model training requires re-conduction for every new lens, limiting the applications of lens-specific CAC methods.
@@
Consequently, developing a universal CAC model trained on datasets under a lens library (LensLib) to address arbitrary optical degradation has hit the forefront. Yet, existing solutions [31, 14] suffer two main limitations: (i) the coverage of the manually collected LensLib is limited and inflexible for expanding, leading to poor zero-shot CAC results when the target lens appears as a new, previously unseen sample, as shown in Fig. 1 (b); (ii) the fine-tuning strategy [31] for model adaptation requires known or calibrated lens descriptions for data preparation, which can hardly be achieved in real-world applications. Recently, the blooming development of Domain Adaptation (DA) [51, 55, 53] provides an effective pipeline to adapt the model trained on sufficient labeled data to unlabeled data with domain shift, which reveals potential in tuning the universal CAC model for a description-unknown lens, but has not been explored.
@@
In this work, we build up a flexible universal CAC framework OmniLens, modeling the CAC as the problem of domain adaptation from a large LensLib to specific lenses. First and foremost, a convincing LensLib covering possible aberration behaviors of real-world lenses is first constructed to pre-train a base model with strong generalization capabilities. To achieve this, instead of manual design or collection, we propose a novel Evolution-based Automatic Optical Design (EAOD) method for automatic LensLib generation, consisting of an evolution mechanism for enriching the diversity of the generated lens structures, and comprehensive optimization objectives to ensure their reasonable aberration behaviors. As shown in Fig. 1 (c), the model pre-trained on automatically generated LensLib generalizes well to the training-unseen lenses. Then, incorporating the easily accessible real-captured images of a specific lens, the LensLib-to-specific domain adaptation is conducted to fine-tune the pre-trained model for better specific CAC results. To this intent, we develop an efficient DA framework based on our statistic observation of Dark Channel Prior (DCP) [37] in optical degradation, which is expressed as an unsupervised regularization term to constrain the CAC model for fast and few-shot domain adaptation. Despite in a blind deconvolution way without access to the lens descriptions, the adapted model can achieve excellent results as shown in Fig. 1 (d).
@@ 2 Related Work
Computational Aberration Correction. Computational Aberration Correction (CAC)  [42, 17] is proposed to enhance the images of low-end lenses with optical degradation. Benefiting from the blooming development of image restoration networks in low-level vision [34, 3, 2], learning-based CAC methods [39, 4, 7] have revealed more impressive results than model-based methods [43, 57, 12]. These methods can be regarded as an implicit modeling of end-to-end non-blind deconvolution, wherein the model learns the deconvolution for a specific lens using data generated from calibrated or simulated lens descriptions. Yet, limited to lens-specific solutions, these models can hardly accommodate new unseen lenses. Recently, universal CAC [31, 14] has been explored for addressing the optical degradation of any lens blindly, where the generalization ability of the model highly depends on the coverage of the LensLib used for training. Existing LensLib construction centered around manually collected lens samples [31, 14], i.e., ZEBASELib, or random Zernike model [20, 23], i.e., ZernikeLib. However, the former suffers from the limited number and coverage, and the latter overlooks the domain shift between generated virtual lenses and real-world ones. Moreover, their fine-tuning pipelines are often not applicable in lens-descriptions-unknown scenarios. Consequently, we propose OmniLens to address these limitations, constructing a more convincing LensLib by automatic optical design, and a LensLib-to-specific domain adaptation pipeline to fine-tune the model with unknown lens descriptions.
@@
Automatic Optical Design. Heuristic algorithms, e.g., Simulated Annealing (SA), Genetic Algorithm (GA), and Particle Swarm Optimization (PSO) have gained widespread adoption in the field of automatic optical design (AOD) [15, 46, 60, 58, 49, 40] for searching the optimized lens structure under the given constraints. Recently, Yang et al. have also proposed to combine curriculum learning and deep optics for achieving better AOD results [56]. However, existing works only focus on finding a certain optimal optical system, ignoring the exploration of possible diverse structures. Efforts have been made to pursue diverse design solutions using Deep Neural Networks (DNN) [10, 11]. Yet, their performance is restricted to the training lens database. Moreover, the lens structures output by the network are not regularized by any stringent quality or physical constraints, which can hardly cover the lenses in practical applications. To this end, we put forward the EAOD method based on heuristic algorithms to generate a diverse range of lenses automatically with machinable structures and reasonable aberration behaviors to cover real-world lenses.
@@
Domain Adaptation in Low-Level Vision. Early effort has been made in mitigating the synthetic-to-real gap problem in CAC via domain adaptation from labeled synthetic domain to the unlabeled target real-world domain [24], which demands a large amount of data and training iterations for learning the domain-mixing representation of optical degradation. Nevertheless, the optical degradation of LensLib varies across each training sample, which can hardly be represented in a uniform way mixed with the target-specific domain. The heavy training cost also renders such approaches unsuitable for the required rapid model fine-tuning in our framework. In addition, there have been works on domain adaption in other low-level vision tasks, e.g., image super-resolution [54, 51], image dehazing [30, 8, 32], and underwater image enhancement [21, 53]. Among them, we find that the utilization of task-specific physical priors (e.g., DCP) [8, 30] contributed to effective adaptation with minimal training cost. To achieve this, we verify the existance of DCP in optical degradation, deriving it into a regularization term to enable the LensLib-to-specific domain adaptation.
@@ 3 Methodology
Figure 3: Overall pipeline of the proposed EAOD method. Taking the lens specifications of 66 piece as an example, the EAOD leverages a hybrid global and local optimization strategy to seek diverse lens structures that maximally satisfy both imaging quality and physical constraints from generation to generation based on an evolution framework with mutation mechanism.
@@ 3.2 Automatic Generation of LensLib
Considering the highly free and empirical process of optical design, designing a large number of lenses manually to traverse possible aberrations of real-world lenses is challenging. To tackle this issue, the idea of AOD [11] is introduced to enable the large-scale generation of lens samples.
@@
The optimization process of $P^{\mathcal{S}^{(j)}}$ is guided by an objective function $\mathcal{L}_{AOD}(P^{\mathcal{S}^{(j)}})$, consisting of both imaging quality constraints and physical structure constraints. Specifically, an imaging quality loss $\mathcal{L}_{IQ}$ calculated by ray-tracing [5] is applied as the main optimization objective, which can be common RMS radius or Modulation Transfer Function (MTF) loss in traditional optical design[45]. Moreover, key physical properties of the lens derived from $P^{\mathcal{S}^{(j)}}$ are constrained by a linear penalty to fall within a reasonable range:
@@
Evolution-based Automatic Optical Design. Based on the Eq. (4), we propose an EAOD method to (i) generate more possible lens structures under one set of specifications, and (ii) ensure the best imaging quality for the current structure for driving the aberration behaviors of the generated samples more close to those of real-world ones.
As shown in Fig. 3, EAOD builds upon a Genetic Algorithm (GA) [26] based evolution framework to produce optimized lens structures from generation to generation. We first define a $Population$ composed of all lens structures to be optimized, where each structure is coined an $Individual$. Without any empirical preference, the $Individual$ is the normalized lens parameters vector $P^{\mathcal{S}^{(j)}}$ from random initialization, and $m$ IndividualsIndividuals are initialized to constitute the $Population$ for enriching diversity:
@@
Then, a hybrid global and local optimization strategy is proposed to find multiple promising high-quality structures $\hat{\mathrm{P}}^{\mathcal{S}^{(j)}}$ for $\mathrm{P}^{\mathcal{S}^{(j)}}$, coined the $Parent$, which minimizes the $\mathcal{L}_{AOD}$. Concretely, the global optimization method, e.g., SA [41] or PSO [9], is employed to rapidly drive the randomly initialized $Individuals$ towards the location of the global optimal solution, where the solutions with superior $\mathcal{L}_{AOD}$ are selected for the following local optimization; while the local optimization method, e.g., ADAM [28] or Damped Least Squares (DLS) [29], aims to seek their local optima for the relatively best imaging results under the primarily optimized structures. In addition, $Parent$ is mutated for more possible structures, where for each IndividualIndividual ${\hat{{P}}^{\mathcal{S}^{(j)}}}_{i}$ in $Parent$, $\gamma$ percentage of its parameters are randomly initialized. The mutated $Parent$ is then mixed with the origin $Parent$ to serve as the start points of the next generation. Finally, the hybrid optimization, selection, and mutation processes are re-conducted for each generation, where the $Parent$ is outputted as the diverse lens structures when the generation meets the set number.
@@ 3.3 Pre-Training on LensLib
We employ a ray-tracing model [5] to calculate the lens descriptions for all samples in ${\mathrm{L}}_{AODLib}$, which are applied to simulate aberration images via patch-wise convolution with clear images and distortion transformation. In this way, the AODLib data pairs $\mathcal{D}_{S}{=}\{{y_{S}}^{(i)},{x}^{(i)}\}_{i=1}^{N}$ are prepared for supervised pre-training using the objective $\mathcal{L}_{CAC}({y^{\prime}_{S}}^{(i)},{x}^{(i)})$, where ${y^{\prime}_{S}}^{(i)}$ is the CAC output. $\mathcal{L}_{CAC}$ can be any training loss in the CAC field, such as the combination of L2 loss and perceptual loss [4], depending on the applied CAC model. The pre-trained universal model can be applied to correct aberrations of real-world lenses in a zero-shot manner. In addition, when the target lens description is available, it can serve as the pre-trained foundation for improving the lens-specific CAC model during few-shot fine-tuning or full model training.
@@ 3.4 LensLib-to-Specific Domain Adaptation
Effective low-level vision DA frameworks [44, 8, 21] are often built upon the statistical priors of degraded images or clean images, e.g., the Dark Channel Prior (DCP) [16] in image dehazing, which can provide an unsupervised regularization for model training. Motivated by the observation from [37] that convolution-induced blur often exhibits DCP, we hypothesize that optical degradation also adheres to DCP, as it originates from the convolution with the PSFs. To validate the hypothesis, we use the DIV2K dataset [50] of natural images, and without loss of generality, choose the second lens in Fig. 1 to construct aberration-clear image pairs. The Dark Channel (DC) image $D(I)$ for an image $I$ can be calculated via a minimum value filtering operation:
@@
where $I^{c}$ is the $c$th color channel, $x,y$ are pixel locations, and $\mathcal{N}(x)$ represents an image patch centered at $x$. We statistically analyze the intensity distribution histograms of the $D(I)$ for all the aforementioned aberration-clear image pairs as shown in Fig. 4, where a set of visual samples of $I$ and $D(I)$ is also provided. Corroborating our hypothesis, the elements in DC images for clear natural images mostly tend towards zero, while aberration images exhibit far more non-zero pixels in their DC images, indicating that optical degradation indeed possesses DCP.

### v3

@@ 1 Introduction
Figure 1: OmniLens is a flexible solution to universal CAC. In real-world cases of unknown lens descriptions for the applied low-end lenses, the pre-trained model with manually collected LensLib [li2021universal] delivers poor performance and can hardly be further fine-tuned. In contrast, the automatically generated LensLib in OmniLens develops a more robust model, where the LensLib-to-specific domain adaptation can further enhance the results.
@@
Computational Aberration Correction (CAC), equipped with a post-image-restoration method to deal with the degradation (i.e., the optical degradation [chen2021optical]) induced by residual optical aberrations of the target optical lens, is a fundamental and long-standing task [schuler2011non, kee2011modeling] in computational imaging. This technology is particularly highly demanded in mobile and wearable vision terminals for lightweight and high-quality photography [suo2023computational, zhang2023large], where low-end lenses [heide2013high, gao2022compact] with simple structure and severe optical degradation are applied, as shown in Fig. 1 (a). Recent advances in CAC have centered around Deep Learning (DL) based methods [lin2022non, chen2021extreme_quality, chen_mobile_2023] equipped with a powerful image restoration network [liang2021swinir, chen2022simple, zamir2022restormer] being trained on the corresponding data pairs under lens-specific optical degradation [chen2021extreme_quality, chen2021optical]. However, these methods have invariably been tailored to a specific lens, revealing limited generalization ability to other different lens designs. In this way, the complex and time-consuming pipeline of lens description calibration, imaging simulation, data preparation, and model training requires re-conduction for every new lens, limiting the applications of lens-specific CAC methods.
@@
Consequently, developing a universal CAC model trained on datasets under a lens library (LensLib) to address arbitrary optical degradation has hit the forefront. Yet, existing solutions [li2021universal, gong2024physics] suffer two main limitations: (i) the coverage of the manually collected LensLib is limited and inflexible for expanding, leading to poor zero-shot CAC results when the target lens appears as a new, previously unseen sample, as shown in Fig. 1 (b); (ii) the fine-tuning strategy [li2021universal] for model adaptation requires known or calibrated lens descriptions for data preparation, which can hardly be achieved in real-world applications. Recently, the blooming development of Domain Adaptation (DA) [wang2021unsupervised, xu2022dual, wang2023domain] provides an effective pipeline to adapt the model trained on sufficient labeled data to unlabeled data with domain shift, which reveals potential in tuning the universal CAC model for a description-unknown lens, but has not been explored.
@@
In this work, we build up a flexible universal CAC framework OmniLens, modeling the CAC as the problem of domain adaptation from a large LensLib to specific lenses. First and foremost, a convincing LensLib covering possible aberration behaviors of real-world lenses is first constructed to pre-train a base model with strong generalization capabilities. To achieve this, instead of manual design or collection, we propose a novel Evolution-based Automatic Optical Design (EAOD) method for automatic LensLib generation, consisting of an evolution mechanism for enriching the diversity of the generated lens structures, and comprehensive optimization objectives to ensure their reasonable aberration behaviors. As shown in Fig. 1 (c), the model pre-trained on automatically generated LensLib generalizes well to the training-unseen lenses. Then, incorporating the easily accessible real-captured images of a specific lens, the LensLib-to-specific domain adaptation is conducted to fine-tune the pre-trained model for better specific CAC results. To this intent, we develop an efficient DA framework based on our statistical observation of Dark Channel Prior (DCP) [pan2016blind] in optical degradation, which is expressed as an unsupervised regularization term to constrain the CAC model for fast and few-shot domain adaptation. Despite being in a blind deconvolution way without access to the lens descriptions, the adapted model can achieve excellent results as shown in Fig. 1 (d).
@@ 2 Related Work
Computational Aberration Correction. Computational Aberration Correction (CAC) [schuler2011non, heide2013high] is proposed to enhance the images of low-end lenses with optical degradation. Benefiting from the blooming development of image restoration networks in low-level vision [liang2021swinir, chen2022simple, chen2022real], learning-based CAC methods [peng2019learned, chen2021extreme_quality, chen_mobile_2023] have revealed more impressive results than model-based methods [Schuler2012Blind, yue2015blind, eboli2022fast]. These methods can be regarded as an implicit modeling of end-to-end non-blind deconvolution, wherein the model learns the deconvolution for a specific lens using data generated from calibrated or simulated lens descriptions. Yet, limited to lens-specific solutions, these models can hardly accommodate new unseen lenses. Recently, universal CAC [li2021universal, gong2024physics] has been explored for addressing the optical degradation of any lens blindly, where the generalization ability of the model highly depends on the coverage of the LensLib used for training. Existing LensLib construction centered around manually collected lens samples [li2021universal, gong2024physics], i.e., ZEBASELib, or random Zernike model [hu2021image, jiang2024computational], i.e., ZernikeLib. However, the former suffers from the limited number and coverage, and the latter overlooks the domain shift between generated virtual lenses and real-world ones. Moreover, their fine-tuning pipelines are often not applicable in lens-descriptions-unknown scenarios. Consequently, we propose OmniLens to address these limitations, constructing a more convincing LensLib by automatic optical design, and a LensLib-to-specific domain adaptation pipeline to fine-tune the model with unknown lens descriptions.
@@
Automatic Optical Design. Heuristic algorithms, e.g., Simulated Annealing (SA), Genetic Algorithm (GA), and Particle Swarm Optimization (PSO) have gained widespread adoption in the field of automatic optical design (AOD) [guo2019new, sun2021automatic, zhang2020automated, yue2022adaptive, tang2019ant, reichert2020development] for searching the optimized lens structure under the given constraints. Recently, Yang et al. have also proposed to combine curriculum learning and deep optics for achieving better AOD results [yang2023curriculum]. However, existing works only focus on finding a certain optimal optical system, ignoring the exploration of possible diverse structures. Efforts have been made to pursue diverse design solutions using Deep Neural Networks (DNN) [cote2019extrapolating, cote2021deep]. Yet, their performance is restricted to the training lens database. Moreover, the lens structures output by the network are not regularized by any stringent quality or physical constraints, which can hardly cover the lenses in practical applications. Heuristic algorithms that consider image quality and physical constraints have recently been used to search diverse initial structures in end-to-end computational imaging systems [gao2025exploring]. The genetic algorithm-based evolution mechanism plays a significant role, which drives overall optimization toward a global optimum while also searching a large number of intermediate local optima. Motivated by this, considering that these intermediate local optima are precisely aberration cases that may occur in real-world lens optimization, we put forward the EAOD method based on heuristic algorithms to generate a diverse range of lenses automatically with machinable structures and reasonable aberration behaviors to cover real-world lenses.
@@
Domain Adaptation in Low-Level Vision. Early effort has been made in mitigating the synthetic-to-real gap problem in CAC via domain adaptation from a labeled synthetic domain to the unlabeled target real-world domain [jiang2025representing], which demands a large amount of data and training iterations for learning the domain-mixing representation of optical degradation. Nevertheless, the optical degradation of LensLib varies across each training sample, which can hardly be represented in a uniform way mixed with the target-specific domain. The heavy training cost also renders such approaches unsuitable for the required rapid model fine-tuning in our framework. In addition, there have been works on domain adaptation in other low-level vision tasks, e.g., image super-resolution [wei2021unsupervised, wang2021unsupervised], image dehazing [li2019semi, chen2021psd, li2023dadrnet], and underwater image enhancement [jiang2022two, wang2023domain]. Among them, we find that the utilization of task-specific physical priors (e.g., DCP) [chen2021psd, li2019semi] contributed to effective adaptation with minimal training cost. To achieve this, we verify the existence of DCP in optical degradation, deriving it into a regularization term to enable the LensLib-to-specific domain adaptation.
@@ 3 Methodology
Figure 3: Overall pipeline of the proposed EAOD method. Taking the lens specifications of 66 piece as an example, the EAOD leverages a hybrid global and local optimization strategy to seek diverse lens structures that maximally satisfy both imaging quality and physical constraints from generation to generation based on an evolution framework with a mutation mechanism.
3.1 Motivation
With recent advances in large datasets and models, image restoration has begun to move toward an all‑in‑one paradigm [jiang2025survey, potlapalli2023promptir, ai2024multimodal]. These developments indicate that training on broadly distributed multiple image degradations enables a universal model to effectively learn many‑to‑one mappings from degradations such as noise, motion blur, rain, fog, and low light to clear images. In addition, in common learning‑based non‑blind CAC pipelines, models are typically trained on aberration data with multiple combinations of perturbations [chen2022computational_mass, jiang2024minimalist] to enhance generalization to unknown manufacturing errors. This strategy effectively prototypes the universal paradigm and shows that training over multiple distributions enables handling diverse aberrations. In this way, this paper explores to construct a large data source covering possible real-world aberration behaviors and pre-train a universal model for aberration correction. Meanwhile, given that generalization is always bounded, we also investigate efficient fine-tuning pipelines based on the pre-trained universal model as a remedy for zero‑shot failure cases, forming a comprehensive and flexible universal aberration correction framework.
@@ 3.3 Automatic Generation of LensLib
Considering the highly free and empirical process of optical design, designing a large number of lenses manually to traverse possible aberrations of real-world lenses is challenging. To tackle this issue, the idea of AOD [cote2021deep] is introduced to enable the large-scale generation of lens samples.
@@
The optimization process of $P^{\mathcal{S}^{(j)}}$ is guided by an objective function $\mathcal{L}_{AOD}(P^{\mathcal{S}^{(j)}})$, consisting of both imaging quality constraints and physical structure constraints. Specifically, an imaging quality loss $\mathcal{L}_{IQ}$ calculated by ray-tracing [chen2021optical] is applied as the main optimization objective, which can be the common RMS radius or Modulation Transfer Function (MTF) loss in traditional optical design[steinich2012optical]. Moreover, key physical properties of the lens derived from $P^{\mathcal{S}^{(j)}}$ are constrained by a linear penalty to fall within a reasonable range:
@@
Evolution-based Automatic Optical Design. Based on the Eq. (4), we propose an EAOD method to (i) generate more possible lens structures under one set of specifications, and (ii) ensure the best imaging quality for the current structure for driving the aberration behaviors of the generated samples closer to those of real-world ones.
As shown in Fig. 3, EAOD builds upon a Genetic Algorithm (GA) [katoch2021review] based evolution framework to produce optimized lens structures from generation to generation. We first define a $Population$ composed of all lens structures to be optimized, where each structure is coined an $Individual$. Without any empirical preference, the $Individual$ is the normalized lens parameters vector $P^{\mathcal{S}^{(j)}}$ from random initialization, and $m$ IndividualsIndividuals are initialized to constitute the $Population$ for enriching diversity:
@@
Then, a hybrid global and local optimization strategy is proposed to find multiple promising high-quality structures $\hat{\mathrm{P}}^{\mathcal{S}^{(j)}}$ for $\mathrm{P}^{\mathcal{S}^{(j)}}$, coined the $Parent$, which minimizes the $\mathcal{L}_{AOD}$. Concretely, the global optimization method, e.g., SA [rutenbar1989simulated] or PSO [985692], is employed to rapidly drive the randomly initialized $Individuals$ towards the location of the global optimal solution, where the solutions with superior $\mathcal{L}_{AOD}$ are selected for the following local optimization; while the local optimization method, e.g., ADAM [kingma2014adam] or Damped Least Squares (DLS) [levenberg1944method], aims to seek their local optima for the relatively best imaging results under the primarily optimized structures. In addition, $Parent$ is mutated for more possible structures, where for each IndividualIndividual ${\hat{{P}}^{\mathcal{S}^{(j)}}}_{i}$ in $Parent$, $\gamma$ percentage of its parameters are randomly initialized. The mutated $Parent$ is then mixed with the origin $Parent$ to serve as the start points of the next generation. Finally, the hybrid optimization, selection, and mutation processes are re-conducted for each generation, where the $Parent$ is outputted as the diverse lens structures when the generation meets the set number.
@@ 3.4 Pre-Training on LensLib
We employ a ray-tracing model [chen2021optical] to calculate the lens descriptions for all samples in ${\mathrm{L}}_{AODLib}$, which are applied to simulate aberration images via patch-wise convolution with clear images and distortion transformation. In this way, the AODLib data pairs $\mathcal{D}_{S}{=}\{{y_{S}}^{(i)},{x}^{(i)}\}_{i=1}^{N}$ are prepared for supervised pre-training using the objective $\mathcal{L}_{CAC}({y^{\prime}_{S}}^{(i)},{x}^{(i)})$, where ${y^{\prime}_{S}}^{(i)}$ is the CAC output. $\mathcal{L}_{CAC}$ can be any training loss in the CAC field, such as the combination of L2 loss and perceptual loss [chen2021extreme_quality], depending on the applied CAC model. The pre-trained universal model can be applied to correct aberrations of real-world lenses in a zero-shot manner. In addition, when the target lens description is available, it can serve as the pre-trained foundation for improving the lens-specific CAC model during few-shot fine-tuning or full model training.
@@ 3.5 LensLib-to-Specific Domain Adaptation
Effective low-level vision DA frameworks [shao2020domain, chen2021psd, jiang2022two] are often built upon the statistical priors of degraded images or clean images, e.g., the Dark Channel Prior (DCP) [he2010single] in image dehazing, which can provide an unsupervised regularization for model training. Motivated by the observation from [pan2016blind] that convolution-induced blur often exhibits DCP, we hypothesize that optical degradation also adheres to DCP, as it originates from the convolution with the PSFs. To validate the hypothesis, we use the DIV2K dataset [timofte2017ntire] of natural images, and without loss of generality, choose the second lens in Fig. 1 to construct aberration-clear image pairs. The Dark Channel (DC) image $D(I)$ for an image $I$ can be calculated via a minimum value filtering operation:
@@
where $I^{c}$ is the $c$th color channel, $x,y$ are pixel locations, and $\mathcal{N}(x)$ represents an image patch centered at $x$. We statistically analyze the intensity distribution histograms of the $D(I)$ for all the aforementioned aberration-clear image pairs as shown in Fig. 4, where a set of visual samples of $I$ and $D(I)$ is also provided. Corroborating our hypothesis, the elements in DC images for clear natural images mostly tend towards zero, while aberration images exhibit far more non-zero pixels in their DC images, indicating that optical degradation indeed possesses DCP (further experimental evidences are provided in Sec. 4.4).